OpenStudy (anonymous):

Evaluate r + s for the given values 1) r= 7 , s= -21 2) r= -23 , s= -12 3) r= 32 , s= -24 4) r= -13 , s= -12 5) r= 35 , s= -15 6) r= -31 , s= 31

OpenStudy (anonymous):

1) 7+-21=14 2) -23+-12=-35 3) 32+-12=20 4) -13+-12=-25 5) 35+-15=20 6) -31+31=0 Before asking, why didn't you use calculator to solve this all? What you had to do is just to put the numbers into r+s and calculate.

OpenStudy (anonymous):

Oh ! Lol and thank yoouu ! But number one is -14 .

OpenStudy (anonymous):

sorry i made typo

OpenStudy (anonymous):

Its ok